OpenStudy (anonymous):

Given that f(x) = x2 − 4x − 3 and g(x) =(x+3)/4solve for f(g(x)) when x = 9. −6 3 6 9

OpenStudy (anonymous):

First g(x) = (9+3)/4 = 3, then plug for f(x) = 3^2 -4(3)-3 = -6 ANS = -6

OpenStudy (anonymous):

ok thank you